OpenStudy (anonymous):

In my book I have \[f(x)=a*(sin(x)+(sin(z)/cos(z))*cos(z))\] And then the next line is \[f(x)=(a/cos(z))*(sin(x)*cos(z)+sin(z)*cos(x))\] Can anyone help me understand what happened :p? Tyvm

OpenStudy (rsadhvika):

multiply first statement with \(\large \frac{\cos z}{\cos z}\)

OpenStudy (anonymous):

Ah yes, lol; that was stupid :p

OpenStudy (rsadhvika):

happens lol when u work too much :)